# Subspace Ordering for Maximum Response Preservation in Sufficient Dimension Reduction

### v1

@@ 5.1 Simulation for Binary Response
We considered parameter configurations from multivariate normal populations with $p=50$ for the MC simulation study. For configurations Q1-Q3 below, we used QDA as the supervised classifier.
@@
For each configuration, we generated 5,000 observations from each population. We varied the training sample sizes to simulate ill-conditioned to well-conditioned estimation of the discriminant parameters. The class-specific training sample sizes for the QDA configurations were $n_{h}=p+1,\ 2p,\ \text{and }5p,\ h\in\{1,2\}$. For the LDA configurations, the training sample sizes were n=p+1, 2p, 5pn=p+1,\ 2p,\ 5p with equal prior class probabilities. We used the remaining observations as the test set. For each SDR method, we projected the training and test sets from $p$ to $d=\text{dim}(\mathcal{S}_{D(Y|\mathbf{X})})$ dimensions, then applied the respective classifier and recorded the estimated conditional error rate, denoted by $\widehat{CER}$. We also recorded the $\widehat{CER}$ of the full-feature data using the respective classifier without dimension reduction. This process was replicated 1,000 times for each configuration. We summarized the MC simulation results in Figures 2 and 3 for the QDA and LDA configurations, respectively, by displaying the distributions of the $\widehat{CER}$s for each SDR method. The horizontal line denotes the median $\widehat{CER}$ of the full-dimensional classifier with no dimension reduction.
From Figure 2, we clearly see that, compared to ordering by eigenvalue magnitude, ordering the DRS by $T_{j}$ can significantly reduce the $\widehat{CER}$. In most cases, we found that eigenvalue ordering often yielded SDR methods with $\widehat{CER}$ values much larger than those of the full-feature QDA, whereas ordering by $T_{j}$ resulted in substantial improvements relative to the full-feature $\widehat{CER}$s. The SIR-II method exhibited the largest gain when it was reordered by the $T_{j}$ criterion. In fact, the $SIR\text{-}II_{T}$ method often achieved the minimum error rates. In contrast, the standard SIR-II consistently produced $\widehat{CER}$s around 0.50. When $n_{h}=p+1=51$ for configuration Q1, the median $\widehat{CER}$ achieved by the full-feature QDA was 0.4287. However, regardless of sample size, the $SIR\text{-}II_{T}$ method yielded an impressive median $\widehat{CER}$ of approximately 0.02. In larger-sample scenarios, the $T_{j}$-ordered SDR methods performed as well as or, in some cases, significantly better than the full-feature QDA. For instance, when $n_{h}=5p=250$ for configuration Q2, the $SAVE_{T}$, $SIR\text{-}II_{T}$, and $SSDR_{T}$ methods achieved the lowest error rates, with $SSDR_{T}$ performing best. Additionally, for configuration Q3 with $n_{h}=250$, the $\textit{PCA}_{T}$ method achieved the minimum median $\widehat{CER}$ of 0.3484, compared to 0.4472 for QDA and 0.4994 for standard PCA. In contrast, all other SDR methods yielded equal or higher error rates than the full-dimensional QDA.
@@ 5.2 Simulation for Continuous Response
The configurations used are similar to those in Zeng et al., (2024). For configurations D1 and D2 below, $\mathbf{X}_{i}$ follows a multivariate normal distribution with $\boldsymbol{\mu}=\mathbf{0}_{p}$ and $\boldsymbol{\Sigma}=AR(0.50)$, where $AR(0.50)$ is a $p\times p$ auto-regressive matrix whose $(i,j)^{\text{th}}$ element is $0.50^{|i-j|}$. For each model, $\varepsilon_{i}$ follows the $\mathcal{N}(0,1)$ distribution. The configurations are as follows:
@@ 6.1 Application to High-Dimensional Gene Expression Data
We attributed the superior performance of the $SIR\text{-}II_{T}$ method to the $T_{j}$ criterion, which provided a data-driven measure that incorporates predictor–response information and yielded a more stable ordering criterion than eigenvalue magnitudes did. In particular, due to singularity issues, the largest eigenvalue was $10^{4}$ and non-unique across the first 7,092 eigenvectors. In contrast, the $T_{j}$ values exhibited clear separation, where the largest value was 5.41, and the second largest was 1.44. The remaining $T_{j}$ values decreased gradually and were unique across all subsequent subspaces. Figures showing the eigenvalue and $T_{j}$ orderings are provided in the Supplementary Material. The eigenvector associated with the largest eigenvalue yielded $T_{1}=0.01$, ranking 6,968th among all $T_{j}$ values whereas the subspace that achieved the largest $T_{j}=5.41$ corresponded to the smallest eigenvalue of 0.0001. Thus, under the traditional eigenvalue-ordering framework, the most informative subspace would not have been selected; however, using the $T_{j}$ criterion, we identified this subspace as the most informative and stable direction.
@@ 6.2 Application to Real Data with a Continuous Response
In this section, we analyzed the Boston housing price data, originally compiled from the 1970 U.S. Census. The dataset contained 506 observations, each representing a census tract within the Boston metropolitan area. For each tract, various socioeconomic, environmental, and structural characteristics were recorded, such as the per capita crime rate by town, average number of rooms per dwelling, full-value property tax rate per \$10,000, and others. The response variable was the median value of owner-occupied homes per \$1,000. See Harrison and Rubinfeld, (1978) for additional details.
Figure 6: Boston housing data discussed in Section 6.2 in the estimated SAVE subspaces corresponding to the largest eigenvalue (left plot) and the largest $F_{j}$ value (right plot).
We randomly partitioned the data into training (80%) and testing (20%) subsets for model estimation and evaluation. As in Section 6.1, we estimated the DRS using the training data for both the SAVE and $SAVE_{F}$ methods, thereby reducing the data dimensionality to $d=1$, which we determined via the training data. Figure 6 presents our results, displaying scatter plots between the median housing prices for the testing set and the reduced testing data in the estimated SAVE and $SAVE_{F}$ subspaces, respectively. From Figure 6, we again found that the subspace associated with the largest eigenvalue captured no response information. In the traditional first SAVE subspace, the resulting testing mean squared error (MSE) from a simple linear regression model was 59.91. The model was clearly insignificant and yielded an MSE comparable to that obtained when we used only the training sample mean, which we found to be 62.82. In contrast, the $F_{j}$ criterion identified a subspace, specifically the fourth SAVE direction, that captured a meaningful predictor-response structure and reduced the testing MSE to 28.05 when it was fitted with a second-degree polynomial model.
This real-data application highlights an important limitation of ordering by eigenvalue magnitude. Although larger eigenvalues typically correspond to directions of greater variability and are often assumed to capture more information, this assumption generally does not hold in practice. In this case, the SAVE subspace associated with the largest eigenvalue did exhibit greater overall variance than the $SAVE_{F}$ subspace. However, this increased variance arose because the eigenvector corresponding to the largest eigenvalue was sparse and effectively projected the data onto two distinct spans. In contrast, the $F_{j}$ criterion considered meaningful variability relative to the response and predictors, which yielded a subspace that captured a more informative structure than just incidental variability.
@@ 7 Discussion
We view this work as an initial step toward developing a broader class of prediction-specific subspace criteria. Our goal is to have demonstrated through the proposed $T_{j}$ and $F_{j}$ criteria that subspace ordering can and often should be informed by the objectives of the supervised learner rather than defaulting to eigenvalue magnitude. In this paper, we considered discrimination and regression settings; however, deriving alternative subspace criteria tailored to other learning objectives remains for future research. Moreover, we assumed that the structural dimension of the CS was known; jointly estimating the structural dimension while simultaneously ordering subspaces using a predictive importance measure is another related future research topic. Such extensions would allow SDR methods to integrate more tightly with the intended supervised model and produce a DRS that is not only statistically sufficient but also optimally informative for prediction.
@@
All reproducible code for the simulations, real data applications, and graphical figures is available at https://github.com/DerikTBoonstra/Subspace_Ordering. Moreover, the $T_{j}$ and $F_{j}$ criteria, along with all of the SDR methods used, are implemented in the working sdr package available at https://github.com/DerikTBoonstra/sdr.

### v2

@@ 5.1 Simulation for Binary Response
We considered parameter configurations from multivariate normal populations that are commonly used for accessing discriminant analysis and similar to those in Mai et al., (2019), Gaynanova and Wang, (2019), and Boonstra et al., (2025). We set $p=50$ for the MC simulation study. For configurations Q1-Q3 below, we used QDA as the supervised classifier.
@@
For each configuration, we generated 5,000 observations from each population. We varied the training sample sizes to simulate ill-conditioned to well-conditioned estimation of the discriminant parameters. The class-specific training sample sizes for the QDA configurations were $n_{h}=p+1,\ 2p,\ \text{and }5p,\ h\in\{1,2\}$. For the LDA configurations, the training sample sizes were n=p+1, 2p,and 5pn=p+1,\ 2p,\ $\text${and }5p with equal prior class probabilities. We used the remaining observations as the test set. For each SDR method, we projected the training and test sets from $p$ to $d=\text{dim}(\mathcal{S}_{D(Y|\mathbf{X})})$ dimensions, then applied the respective classifier and recorded the estimated conditional error rate, denoted by $\widehat{CER}$. We also recorded the $\widehat{CER}$ of the full-feature data using the respective classifier without dimension reduction. This process was replicated 1,000 times for each configuration. We summarized the MC simulation results in Figures 2 and 3 for the QDA and LDA configurations, respectively, by displaying the distributions of the $\widehat{CER}$s for each SDR method. The horizontal line denotes the median $\widehat{CER}$ of the full-dimensional classifier with no dimension reduction.
@@
From Figure 2, we clearly see that, compared to ordering by eigenvalue magnitude, ordering the DRS by $T_{j}$ can significantly reduce the $\widehat{CER}$. In most cases, we found that eigenvalue ordering often yielded SDR methods with $\widehat{CER}$ values much larger than those of the full-feature QDA, whereas ordering by $T_{j}$ resulted in substantial improvements relative to the full-feature $\widehat{CER}$s. The SIR-II method exhibited the largest gain when it was reordered by the $T_{j}$ criterion. In fact, the $SIR\text{-}II_{T}$ method often achieved the minimum error rates. In contrast, the standard SIR-II consistently produced $\widehat{CER}$s around 0.50. When $n_{h}=p+1=51$ for configuration Q1, the median $\widehat{CER}$ achieved by the full-feature QDA was 0.4287. However, regardless of sample size, the $SIR\text{-}II_{T}$ method yielded an impressive median $\widehat{CER}$ of approximately 0.02. In larger-sample scenarios, the $T_{j}$-ordered SDR methods performed as well as or, in some cases, significantly better than the full-feature QDA. For instance, when $n_{h}=5p=250$ for configuration Q2, the $SAVE_{T}$, $SIR\text{-}II_{T}$, and $SSDR_{T}$ methods achieved the lowest error rates, with $SSDR_{T}$ performing best. For configuration Q3 with $n_{h}=250$, the $\textit{PCA}_{T}$ method achieved the minimum median $\widehat{CER}$ of 0.3484, compared to 0.4472 for QDA and 0.4994 for standard PCA. In contrast, all other SDR methods yielded equal or higher error rates than the full-dimensional QDA.
@@ 5.2 Simulation for Continuous Response
The configurations used are similar to those in Lin et al., (2019) and Zeng et al., (2024). For configurations D1 and D2 below, $\mathbf{X}_{i}$ follows a multivariate normal distribution with $\boldsymbol{\mu}=\mathbf{0}_{p}$ and $\boldsymbol{\Sigma}=AR(0.50)$, where $AR(0.50)$ is a $p\times p$ auto-regressive matrix whose $(i,j)^{\text{th}}$ element is $0.50^{|i-j|}$. For each model, $\varepsilon_{i}$ follows the $\mathcal{N}(0,1)$ distribution. The configurations are as follows.
@@ 6.1 Application to High-Dimensional Gene Expression Data
We attributed the superior performance of the $SIR\text{-}II_{T}$ method to the $T_{j}$ criterion, which provided a data-driven measure that yielded a more stable ordering criterion than eigenvalue magnitude did. In particular, due to singularity issues, the largest eigenvalue was $10^{4}$ and non-unique across the first 7,092 eigenvectors. In contrast, the $T_{j}$ values exhibited clear separation, where the largest value was 5.41, and the second largest was 1.44. The remaining $T_{j}$ values decreased gradually and were unique across all subsequent subspaces. Figures showing the eigenvalue and $T_{j}$ orderings are provided in the Supplementary Material. The eigenvector associated with the largest eigenvalue yielded $T_{1}=0.01$, ranking 6,968th among all $T_{j}$ values whereas the subspace that achieved the largest $T_{j}=5.41$ corresponded to the smallest eigenvalue of 0.0001. Thus, under the traditional eigenvalue-ordering framework, the most informative subspace would not have been selected; however, using the $T_{j}$ criterion, we identified this subspace as the most informative and stable direction.
@@ 6.2 Application to Real Data with a Continuous Response
In this section, we analyzed the Boston housing price data, originally compiled from the 1970 U.S. Census. The dataset contained 506 observations, each representing a census tract within the Boston metropolitan area. For each tract, 13 various economic and environmental characteristics were recorded, such as the per capita crime rate by town, average number of rooms per dwelling, full-value property tax rate per \$10,000, and others. The response variable was the median value of owner-occupied homes per \$1,000. See Harrison and Rubinfeld, (1978) for additional details. Consistent with previous studies (e.g., Chen et al., (2010)), we removed the observations with crime rates larger than 3.2 because they coincided with constant predictors. Thus, the data used for analysis had 374 observations.
Figure 6: Boston housing data discussed in Section 6.2 in the estimated PCA subspaces corresponding to the largest eigenvalue (left plot) and the largest $F_{j}$ value (right plot). Estimated simple ordinary least squares solution using the training data is given by the line.
We randomly partitioned the data into training (80%) and testing (20%) subsets for model estimation and evaluation. As in Section 6.1, we estimated the DRS using the training data for both the PCA and $PCA_{F}$ methods, thereby reducing the data dimensionality to $d=1$, which we determined via the training data. Figure 6 presents our results, displaying scatter plots between the median housing prices for the testing set and the reduced testing data in the estimated PCA and $PCA_{F}$ subspaces, respectively. From Figure 6, we again found that the subspace associated with the largest eigenvalue captured very little response information. In the traditional first PCA subspace, the resulting testing mean squared error (MSE) from a simple linear regression model was 67.64. In contrast, the $F_{j}$ criterion identified a subspace, specifically the 11th PCA direction, that captured a meaningful predictor-response structure and reduced the testing MSE to 22.12.
This real-data application highlights an important limitation of ordering by eigenvalue magnitude. Although larger eigenvalues typically correspond to directions with greater variability in the data and are often assumed to capture more information, this assumption generally does not hold in practice. In this application, the data did in fact exhibit greater overall variability in the PCA subspace associated with the largest eigenvalue than in the first $PCA_{F}$ subspace. However, this increased variability arose because the eigenvector corresponding to the largest eigenvalue was sparse and effectively projected the data onto two distinct spans. In contrast, the $F_{j}$ criterion considered meaningful variability relative to the response and predictors, which yielded a subspace that captured a more informative structure than just incidental variance.
@@ 7 Discussion
We view this work as an initial step toward developing a broader class of prediction-specific subspace ordering criteria. Our goal is to have demonstrated through the proposed $T_{j}$ and $F_{j}$ criteria that subspace ordering can and often should be informed by the objectives of the supervised learner rather than defaulting to eigenvalue magnitude. In this paper, we considered discrimination and regression settings; however, deriving alternative subspace criteria tailored to other learning objectives remains for future research. Moreover, we assumed that the structural dimension of the CS was known; jointly estimating the structural dimension while simultaneously ordering subspaces using a predictive importance measure is another related future research topic. Such extensions would allow SDR methods to integrate more tightly with the intended supervised model and produce a DRS that is not only statistically sufficient but also optimally informative for prediction.
@@
All reproducible code for the simulations, real data applications, and graphical figures is available at https://github.com/DerikTBoonstra/Subspace_Ordering . Moreover, the $T_{j}$ and $F_{j}$ criteria, along with all of the SDR methods used, are implemented in the working sdr R package available at https://github.com/DerikTBoonstra/sdr .